A shopper does 110.0 J of work in pushing cart by a force of 33.0 N to a distance of 8.25 m. At what angle is the force applied with respect to the ground
1
Expert's answer
2018-08-20T10:09:08-0400
The formula for work is
W=Fd cos⁡α
cos⁡α=W/Fd
α=cos^(-1)⁡〖W/Fd〗=cos^(-1)⁡〖110/33(8.25) 〗=66.2°.